Provide a reasonable pathway for the making of CH4 from Na2CO3
Answer #1

From Na2CO3:

Na2CO3 + 2 HCl = 2 NaCl + CO2 + H2O

CO2 is treated:

CO2 + H2 = CH4 + H2O

If you liked the answer, please rate it in a positive way, you would help me a lot, thank you.